A beam of monochromatic blue light of wavelength 4200 Å in air travels in water of refractive index 4/3. Its wavelength in water
will be:
(A) 4200 Å (B) 5800 Å (C) 4150 Å (D) 3150 Å

Grade:

2 Answers

Shubham paliwal
5 years ago
Ans. Let wavelength in water = x
         So,
        4/3 = 4200/x 
        
        4x= 4200*3
         
        Therfore,
         x = 4200*3/4
         x = 12600/4
        So, x = 3150 (D)
MATISH BAL
5 years ago
4/3 =3x10^8/v  ,v=2.25x10^8......(1) ,c=frequency x wavelength ,frequency= 7.14x10^14 hz....(2) ,comparing (1) and(2) ,we get,velocity=frequency x wavelength ,therefore, wavelength= (1)/(2) =3150 A